OpenStudy (anonymous):

How do I simplify 3 sqrt 6 +5 / sqrt 6 + 2

OpenStudy (anonymous):

is this correct? \[3\sqrt{6} + 5 / \sqrt{6} +2\]

OpenStudy (anonymous):

am i reading it write is what i mean?

OpenStudy (anonymous):

yes. how do i simplify that?

OpenStudy (anonymous):

ok, you want to simplify the denominator, do this by multiplying the top and bottom of the equation by the conjugate of the denominator (\[\sqrt{6} - 2\])

OpenStudy (luigi0210):

Hello @whatwhatt Welcome to Openstudy :)

OpenStudy (anonymous):

this will leave you with \[(3\sqrt{6}+5)(\sqrt{6}-2)\]

OpenStudy (anonymous):

all divided by 2

OpenStudy (anonymous):

why is it divided by 2?

OpenStudy (anonymous):

\[\frac{ (3\sqrt{6}+5)(\sqrt{6}−2) }{ 2 }\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then you just need to expand the brackets in the numerator

OpenStudy (anonymous):

do you know how to do that or did you want me to step you through?

OpenStudy (anonymous):

would you please take me through that too?

OpenStudy (anonymous):

yeah sure, lemme write it out in my book real quick so i can explain it as clear as i can

OpenStudy (anonymous):

i dont really understanding multiplying with radicals

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

did you get it @Omistone

OpenStudy (anonymous):

\[=\frac{ (3\sqrt{6}+5)(\sqrt{6}-2) }{ 2 }\] \[= \frac{ 18-6\sqrt{6}+5\sqrt{6}-10 }{ 2 }\] \[= \frac{18-\sqrt{6}-10 }{ 2 } \] \[=9-\frac{ \sqrt{6} }{ 2 }-5\] which = 2.775255129

OpenStudy (anonymous):

sorry, that took a really long time to write in the equation writer

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

no problem :) always glad to help with surds